# ACE-RTL: When Agentic Context Evolution Meets RTL-Specialized LLMs

### v1

@@ Abstract.
Recent advances in large language models (LLMs) have sparked growing interest in applying them to hardware design automation, particularly for accurate RTL code generation. Prior efforts follow two largely independent paths: (i) training domain-adapted RTL models to internalize hardware semantics, (ii) developing agentic systems that leverage frontier generic LLMs guided by simulation feedback. However, these two paths exhibit complementary strengths and weaknesses. In this work, we present ACE-RTL that unifies both directions through Agentic Context Evolution (ACE). ACE-RTL integrates an RTL-specialized LLM, trained on a large-scale dataset of 1.7 million RTL samples, with a frontier reasoning LLM through three synergistic components: the generator, reflector, and coordinator. These components iteratively refine RTL code toward functional correctness. We further introduce a parallel scaling strategy that significantly reduces the number of iterations required to reach correct solutions. On the Comprehensive Verilog Design Problems (CVDP) benchmark, ACE-RTL achieves up to a $44.87\%$ pass rate improvement over $14$ competitive baselines while requiring only four iterations on average.
@@
copyright: none11footnotetext: Work done during affiliation with NVIDIA, now at Agentrys.
1. Introduction
@@
Large language models (LLMs) have shown rapid progress in software development and are increasingly explored for hardware design automation (Jimenez et al., 2023; Cursor AI, 2025; Anthropic, 2025b; Liu et al., 2023a; Deng et al., 2025a). As hardware systems continue to grow in scale and complexity, developing register-transfer-level (RTL) designs remains a labor-intensive and error-prone process requiring deep domain expertise. Leveraging LLMs to translate natural-language intent into functionally correct RTL code can significantly accelerate design cycles. Moreover, accurate RTL generation is essential for downstream workflows such as performance optimization and architectural design space exploration (Zhang et al., 2025; Yao et al., 2024; Wang et al., 2025).
@@
In parallel to this motivation, the research community has pursued two main directions, as shown in Figures 1(a) and 1(b), for applying LLMs to RTL code generation. (i) The first focuses on developing domain-specialized LLMs trained on RTL data, where models are fine-tuned to internalize RTL syntax and hardware design semantics. For instance, RTLCoder (Liu et al., 2024c, 2025) develops a specialized model fine-tuned on a dedicated Verilog dataset. CraftRTL (Liu et al., 2024b) further enhances RTL coding capability by training on synthetic data that capture non-textual design representations. More recently, ScaleRTL (Deng et al., 2025b) scales up post-training on a large RTL reasoning corpus, yielding a strong specialized model that achieves state-of-the-art (SoTA) results on mainstream RTL benchmarks. (ii) The second direction centers on agentic systems that leverage frontier generic LLMs together with RTL-specific tools such as simulators and waveform analyzers. Along this line, VerilogCoder (Ho et al., 2025) introduces a multi-agent framework based on GPT4-Turbo (OpenAI, 2024) that decomposes RTL design into fine-grained sub-tasks through a task-and-circuit relation graph and repairs functional errors using a waveform tracing tool. Similarly, MAGE (Zhao et al., 2024) employs multiple Claude3.5-Sonnet (Anthropic, 2024) LLM agents tightly integrated with hardware-domain toolchains to produce functionally correct RTL code.
@@
Despite these advances, the two research directions remain largely orthogonal and exhibit complementary limitations. Training-based approaches often yield RTL-specialized models with limited general capabilities such as long-context reasoning, multi-step planning, and instruction following, reducing their effectiveness on complex design tasks. Conversely, agent-based methods lack RTL-specific knowledge learned from large-scale hardware datasets, causing the system to struggle on problems that require deep hardware semantics and domain expertise. These limitations highlight the need for a unified framework that integrates an RTL-specialized model with a frontier general-purpose LLM, enabling the system to leverage the strengths of both paradigms.
Refer to caption Placeholder.
Figure 1. Comparison of prior works and ACE-RTL.
@@
In this work, we advance LLM-based RTL coding by synergizing the complementary strengths of an RTL-specialized model and a frontier general-purpose LLM. To this end, we propose ACE-RTL, which adopts the notion of Agentic Context Evolution (ACE) through three components: the Generator, the Reflector, and the Coordinator. As illustrated in Figure 2(a), the Generator is an RTL-specialized LLM trained on 1.7 million diverse RTL samples. When the Generator produces buggy code, the Reflector, powered by the frontier LLM Claude4-Sonnet (Anthropic, 2025a), is invoked to analyze simulation feedback, trace errors back to faulty logic, and provide high-level fix guidance. The Coordinator then incorporates this guidance into an incrementally evolving context that guides the Generator in the next iteration of code generation. By iteratively cycling among the Generator, Reflector, and Coordinator, ACE-RTL effectively combines domain-specific RTL expertise with strong general-purpose reasoning to produce functionally correct RTL designs. We further introduce a parallel scaling strategy that launches multiple ACE-RTL processes simultaneously and terminates the remaining ones once any process succeeds. This significantly reduces the total number of iterations needed to obtain correct RTL implementations.
@@
To evaluate the efficacy of ACE-RTL, we compare it against the strongest frontier LLMs, as well as SoTA RTL-specialized models. More importantly, instead of relying on prior RTL benchmarks such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2024) and RTLLM (Lu et al., 2024), which primarily consist of relatively simple specification-to-RTL generation tasks (e.g., small modules or textbook-style problems), we conduct evaluations on the Comprehensive Verilog Design Problems (CVDP) benchmark (Pinckney et al., 2025). CVDP encompasses hundreds of diverse and challenging RTL coding tasks, including code completion, specification-to-RTL generation, code modification, and debugging. Since earlier benchmarks have become saturated and no longer sufficiently differentiate recent approaches, we argue that evaluating on CVDP provides a more rigorous and practical assessment of model capability and more convincingly demonstrates meaningful progress in this domain. We summarize our technical contributions as follows:
$\bullet$ To our knowledge, this work is the first to introduce the ACE framework for RTL coding, which jointly leverages a frontier generic LLM and an RTL-specialized model.
$\bullet$ Motivated by data scaling laws, we construct a 1.7M-sample RTL dataset covering diverse coding problems. Trained on this corpus, our RTL-specialized model surpasses SoTA baselines, including GPT-5 and ScaleRTL, by up to $21.45\%$.
@@
$\bullet$ By integrating the RTL-specialized model with the ACE framework, our proposed ACE-RTL achieves up to a $44.87\%$ improvement in pass rate over all baselines on the CVDP benchmark, demonstrating the effectiveness of ACE for complex RTL tasks.
@@
$\bullet$ With parallel scaling, ACE-RTL requires only four iterations on average to produce correct RTL code, achieving up to a $3.06\times$ reduction in total iterations over the non-scaled version.
2. Background
In this section, we review prior studies on RTL-specialized LLMs, agentic systems, and RTL benchmarking, which together establish the foundation and motivation for ACE-RTL.
2.1. Domain-Adapted LLMs for RTL Tasks
@@
Adapting LLMs for RTL coding via model training has attracted growing attention due to their potential to improve hardware design productivity. RTLCoder (Liu et al., 2024c, 2025) develops a GPT-assisted pipeline for automated RTL data synthesis, while CraftRTL (Liu et al., 2024b) generates synthetic, correct-by-construction samples derived from design artifacts. OriGen (Cui et al., 2024b) adopts compiler-guided reflection and augmentation to improve functional correctness. Recently, ScaleRTL (Deng et al., 2025b) scales both training data and test-time compute by fine-tuning on billions of RTL reasoning tokens and applying iterative reasoning extension, achieving SoTA results on VerilogEval (Pinckney et al., 2024) and RTLLM (Lu et al., 2024) benchmarks. These efforts collectively advance the development of RTL-specialized models. However, it remains unclear how to effectively integrate such models into an agentic system that can further push the frontier of LLM-based RTL coding.
2.2. Agentic Systems and Context Evolution
@@
Recent agentic systems such as VerilogCoder (Ho et al., 2025) and MAGE (Zhao et al., 2024) have demonstrated the potential of multi-agent collaboration for RTL code generation. VerilogCoder decomposes RTL design into fine-grained subtasks using a Task and Circuit Relation Graph (TCRG) and employs a waveform-tracing debugger to locate functional errors, while MAGE leverages multiple specialized agents for RTL generation, testbench creation, and debugging through simulator-based feedback. Despite their promising results, both methods depend heavily on complex, tool-specific planning and lack an RTL-specialized model that captures broad hardware coding logic, which makes them difficult to generalize beyond predefined task structures. In this work, we take a different perspective: we argue that the key role of an agentic system is to dynamically construct the right context to guide the RTL generator toward correct solutions. To this end, ACE-RTL iteratively evolves contextual information such as prior attempts and feedback to progressively steer an RTL-specialized model toward functional correctness.
2.3. RTL Benchmarks for LLM Evaluation
@@
Prior mainstream benchmarks such as VerilogEval and RTLLM mostly evaluate small, self-contained RTL problems, emphasizing short specification-to-RTL snippets rather than sustained reasoning over realistic design contexts. RTL-Repo (Allam and Shalan, 2024) takes a step toward practical settings by introducing GitHub-derived projects and prompting models to complete redacted regions, yet it remains limited to code completion and omits broader challenges such as specification-to-RTL generation, code modification, or debugging. We argue that rigorous assessment requires benchmarks stressing both context length and task diversity. Therefore, this work focuses on evaluating RTL methods on CVDP (Pinckney et al., 2025), whose problem statements and target code are several orders of magnitude longer than those in VerilogEval. CVDP includes hundreds of complex RTL problems spanning multiple tasks and hardware domains such as arithmetic units, on-chip protocols, memory hierarchies, DSP kernels, and communication blocks. Empirically, we observe that CVDP better distinguishes model capability and more faithfully reflects the demands of practical RTL design.
3. The ACE-RTL Framework
Refer to caption Placeholder.
Figure 2. Overview of ACE-RTL and its parallel scaling strategy.
@@ 3. The ACE-RTL Framework
Figure 2 provides an overview of ACE-RTL, which integrates an RTL-specialized model and a frontier generic LLM through three synergistic components: the Generator, Reflector, and Coordinator. At a high level, ACE-RTL operates in an iterative agentic loop that progressively refines RTL code toward functional correctness. These three components are discussed in detail in Sections 3.1, 3.2, and 3.3, respectively. In addition, Section 3.4 introduces a parallel scaling strategy that executes multiple ACE-RTL instances concurrently to speedup convergence efficiency while preserving correctness.
3.1. Generator: An RTL-Specialized LLM
We train an RTL-specialized model as ACE-RTL-Generator, which aims to produce correct code, conditioned on an input specification and an evolving context constructed during prior iterations. To this end, we follow two steps to construct a large-scale RTL dataset: (i) collecting and filtering raw RTL scripts from public sources, (ii) generating diverse specification–code pairs via in-context learning.
High-Quality RTL Script Collection
Motivated by data scaling laws in LLMs (Hoffmann et al., 2022; Kaplan et al., 2020), we begin by collecting five million raw RTL scripts from public repositories and open hardware projects. To ensure data quality, we perform a sequence of filtering and validation steps. First, we remove duplicate files and eliminate machine-generated code, including netlist scripts or RTL generated from high-level synthesis. We also filter out extreme outliers by excluding scripts with fewer than $30$ lines or more than $2,000$ lines. Next, we apply syntax validation using the Icarus Verilog (iverilog) (Williams and Baxter, 2002) to discard all code samples containing syntax errors.
To prevent contamination of downstream evaluation benchmarks, we use the Jaccard similarity metric defined in (Deng et al., 2025b) to measure the similarity between each collected RTL snippet and all benchmark golden solutions. Any collected code whose similarity score exceeds $0.8$ with respect to any golden solution is discarded. This effectively prevents overlap with test data and ensures fair evaluation. After all filtering stages, we retain 157K high-quality RTL scripts as the foundation for specification-code pair generation.
Specification and Code Pair Generation
Given the filtered RTL corpus, we employ in-context learning to automatically generate diverse specifications for each RTL design. We first manually construct a pool of 32 high-quality examples, each containing a specification paired with its RTL implementation. These examples span multiple tasks, including RTL code generation, modification, and debugging. For each raw RTL script, we randomly sample one instance from this pool as context and prompt multiple LLMs, including open-source models such as GPT-OSS-120B (Agarwal et al., 2025) and DeepSeek-R1 (Guo et al., 2025) as well as proprietary ones, to produce diverse pairs of task-specific specifications and corresponding golden RTL implementations. Since the generated RTL may deviate from the original design, we apply an additional filtering stage to remove samples with syntax errors or potential benchmark contamination, identified by high similarity to downstream evaluation benchmarks. For modification and debugging tasks, the specification includes a simplified or intentionally faulty version of the golden code, emulating realistic design correction workflows. Through this pipeline, we curate a large-scale dataset of 1.7 million specification–RTL pairs that comprehensively cover diverse hardware design scenarios.

### v2

@@ Abstract
Recent advances in LLMs have sparked growing interest in applying them to hardware design automation, particularly for accurate RTL code generation. Prior efforts follow two largely independent paths: (i) training domain-adapted RTL models to internalize hardware semantics, (ii) developing agentic systems that leverage frontier generic LLMs guided by simulation feedback. However, these two paths exhibit complementary strengths and weaknesses. In this work, we present ACE-RTL that unifies both directions through Agentic Context Evolution (ACE). ACE-RTL integrates an RTL-specialized LLM, trained on a large-scale dataset of 1.7 million RTL samples, with a frontier reasoning LLM through three synergistic components: the generator, reflector, and coordinator. These components iteratively refine RTL code toward functional correctness. We further analyze a parallel scaling strategy that reduces wall-clock iterations to first success by exploring diverse debugging trajectories concurrently. On the CVDP benchmark, ACE-RTL achieves up to a $41.02\%$ pass rate improvement over $14$ competitive baselines.
@@
11footnotetext: Work done during affiliation with NVIDIA, now at Agentrys.
I Introduction
@@
Large language models (LLMs) have shown rapid progress in software development and are increasingly explored for hardware design automation [16, 8, 5, 19, 9]. As hardware systems continue to grow in scale and complexity, developing register-transfer-level (RTL) designs remains a labor-intensive and error-prone process requiring deep domain expertise. Leveraging LLMs to translate natural-language intent into functionally correct RTL code can significantly accelerate design cycles. Moreover, accurate RTL generation is essential for downstream workflows such as performance optimization and architectural design space exploration [35, 34, 30].
@@
In parallel to this motivation, the research community has pursued two main directions, as shown in Figures 1(a) and 1(b), for applying LLMs to RTL code generation. (i) The first focuses on developing domain-specialized LLMs trained on RTL data, where models are fine-tuned to internalize RTL syntax and hardware design semantics. For instance, RTLCoder [23, 22] develops a specialized model fine-tuned on a dedicated Verilog dataset. CraftRTL [21] further enhances RTL coding capability by training on synthetic data that capture non-textual design representations. More recently, ScaleRTL [10] scales up post-training on a large RTL reasoning corpus, yielding a strong specialized model that achieves state-of-the-art (SoTA) results on mainstream RTL benchmarks. (ii) The second direction centers on agentic systems that leverage frontier generic LLMs together with RTL-specific tools such as simulators and waveform analyzers. Along this line, VerilogCoder [13] introduces a multi-agent framework based on GPT4-Turbo [25] that decomposes RTL design into fine-grained sub-tasks through a task-and-circuit relation graph and repairs functional errors using a waveform tracing tool. Similarly, MAGE [37] employs multiple Claude [3] LLM agents tightly integrated with hardware-domain toolchains to produce functionally correct RTL code.
@@
Despite these advances, the two research directions remain largely orthogonal and exhibit complementary limitations. Training-based approaches often yield RTL-specialized models with limited general capabilities such as long-context reasoning, multi-step planning, and instruction following, reducing their efficacy on complex design tasks. Conversely, agent-based methods lack hardware knowledge learned from large-scale RTL datasets, causing the system to struggle on problems that require deep hardware semantics and domain expertise. These limitations highlight the need for a unified framework that integrates an RTL-specialized model with a frontier general-purpose LLM, enabling the system to leverage the strengths of both paradigms.
Figure 1: Comparison of prior works and ACE-RTL.
@@
In this work, we advance LLM-based RTL coding by synergizing the complementary strengths of an RTL-specialized model and a frontier general-purpose LLM. To this end, we propose ACE-RTL, which adopts the notion of Agentic Context Evolution (ACE) through three components: the Generator, the Reflector, and the Coordinator. As illustrated in Figure 2(a), the Generator is an RTL-specialized LLM trained on 1.7 million diverse RTL samples. Different from prior RTL models that are trained exclusively on specification-to-RTL data, our dataset additionally includes code editing and debugging tasks. This design better prepares the Generator for its role in the agentic loop, where it not only generates initial RTL code from specification, but also edits and fixes bugs in code in later iterations. When errors occur, the Reflector, powered by the frontier LLM Claude4-Sonnet [4], analyzes simulation feedback, traces errors to faulty logic, and provides high-level fix guidance. The Coordinator then incorporates this guidance into an incrementally evolving context that guides the Generator in the next iteration of code generation. By iteratively cycling among the Generator, Reflector, and Coordinator, ACE-RTL effectively combines domain-specific RTL expertise with strong general-purpose reasoning to produce functionally correct RTL designs. We further introduce a parallel scaling strategy that launches multiple ACE-RTL processes simultaneously and terminates the remaining ones once any process succeeds. This largely reduces the iterations required to reach a correct RTL implementation, with only a modest increase in token usage.
@@
To evaluate the efficacy of ACE-RTL, we compare it against the strongest frontier LLMs, as well as SoTA RTL-specialized models and agents. More importantly, instead of relying on prior RTL benchmarks such as VerilogEval [20, 27] and RTLLM [24], which primarily consist of relatively simple specification-to-RTL generation tasks (e.g., small modules or textbook-style problems), we conduct evaluations on the Comprehensive Verilog Design Problems (CVDP) benchmark [28]. CVDP encompasses hundreds of diverse and challenging RTL coding tasks. Since earlier benchmarks have become saturated and no longer sufficiently differentiate recent approaches, we argue that evaluating on CVDP provides a more rigorous and practical assessment of model capability and more convincingly demonstrates meaningful progress in this domain. We summarize our technical contributions as follows:
$\bullet$ To our knowledge, ACE-RTL is the first approach to integrate an RTL-specialized model into an agentic system through iterative context evolution, demonstrating that a finetuned RTL model can contribute not only as a strong standalone generator, but also as an effective core component in iterative agentic refinement.
$\bullet$ Trained on a 1.7M-sample RTL corpus spanning diverse tasks, our RTL model achieves up to a $21.45\%$ higher pass rate than strong standalone baselines such as Claude4 and GPT5, while also aligning well with the iterative bug-fixing setting used in our agentic loop.
@@
$\bullet$ By combining this RTL model with the ACE framework, ACE-RTL achieves up to a $41.02\%$ pass-rate improvement over all baselines on the CVDP benchmark, demonstrating the effectiveness of iterative context evolution for complex RTL tasks.
@@
$\bullet$ With parallel scaling, ACE-RTL speeds up convergence to correct RTL code, achieving a $2.77\times$ reduction in wall-clock iterations relative to the non-scaled version, with only a modest increase in token usage.
II Background
II-A Domain-Adapted LLMs for RTL Tasks
@@
Adapting LLMs for RTL coding via model training has attracted growing attention due to their potential to improve hardware design productivity. RTLCoder [23, 22] develops a GPT-assisted pipeline for automated RTL data synthesis, while CraftRTL [21] generates synthetic, correct-by-construction samples derived from design artifacts. OriGen [7] adopts compiler-guided reflection and augmentation to improve functional correctness. Recently, ScaleRTL [10] scales both training data and test-time compute by fine-tuning on billions of RTL reasoning tokens and applying iterative reasoning extension, achieving SoTA results on VerilogEval [27] and RTLLM [24] benchmarks. These efforts collectively advance the development of RTL-specialized models. However, those prior models are trained predominantly on specification-to-RTL generation data, with limited coverage of code editing and debugging behaviors. This training mismatch makes them less effective in practical agentic settings, where the model needs to iteratively repair buggy RTL produced in earlier iterations. To tackle this issue, we curate a 1.7M-sample RTL corpus covering code generation, editing, and debugging tasks. Training on this corpus yields an RTL-specialized model that is not only strong at standalone RTL generation, but also fits agentic workflows for iterative bug fixing.
II-B Agentic Systems and Context Evolution
@@
Recent agentic systems such as VerilogCoder [13] and MAGE [37] have demonstrated the potential of multi-agent collaboration for RTL code generation. VerilogCoder decomposes RTL design into fine-grained subtasks using a Task and Circuit Relation Graph (TCRG) and employs a waveform-tracing debugger to locate functional errors, while MAGE leverages multiple specialized agents for RTL generation, testbench creation, and debugging through simulator-based feedback. Despite their promising results, both methods depend heavily on complex, tool-specific planning and lack an RTL-specialized model that captures broad hardware coding logic, which can make them harder to generalize beyond predefined task structures. In this work, we take a different perspective: we argue that the key role of an agentic system is to dynamically construct the right context to guide the RTL generator toward correct solutions. To this end, ACE-RTL iteratively evolves contextual information from previous iterations to progressively steer an RTL model toward functional correctness.
II-C RTL Benchmarks for LLM Evaluation
@@
Prior mainstream benchmarks such as VerilogEval and RTLLM mostly evaluate small, self-contained RTL problems, emphasizing short specification-to-RTL snippets rather than sustained reasoning over realistic design contexts. RTL-Repo [2] takes a step toward practical settings by introducing GitHub-derived projects and prompting models to complete redacted regions, yet it remains limited to code completion and omits broader challenges such as specification-to-RTL generation, code modification, or debugging. We argue that rigorous assessment requires benchmarks stressing both context length and task diversity. Therefore, this work focuses on evaluating RTL methods on CVDP [28], whose problem statements and target code are several orders of magnitude longer than those in VerilogEval. CVDP includes hundreds of complex RTL problems spanning multiple tasks and hardware domains such as arithmetic units, on-chip protocols, memory hierarchies, DSP kernels, and communication blocks. Empirically, we observe that CVDP better distinguishes model capability and more faithfully reflects the demands of practical RTL design.
III The ACE-RTL Framework
Figure 2: Overview of ACE-RTL and its parallel scaling strategy.
@@ III The ACE-RTL Framework
Figure 2 provides an overview of ACE-RTL, which integrates an RTL-specialized model and a frontier generic LLM through three synergistic components: the Generator, Reflector, and Coordinator. At a high level, ACE-RTL operates in an iterative agentic loop that progressively refines RTL code toward functional correctness. These three components are discussed in detail in Sections III-A, III-B, and III-C, respectively. In addition, Section III-D introduces a parallel scaling strategy that executes multiple ACE-RTL instances concurrently to speedup convergence efficiency.
III-A Generator: An RTL-Specialized LLM
We develop an RTL-specialized model, referred to as ACE-RTL-Generator, that is trained on a large-scale synthetic dataset to capture diverse hardware design patterns and semantics. Unlike prior synthetic data generation (SDG) approaches that focus primarily on specification-to-RTL translation, our SDG considers a broader family of tasks that better reflect real-world hardware development workflows, including code editing and debugging. These tasks enable the model to learn refinement and error-correction capabilities that are essential for RTL design, particularly when deployed within an agentic workflow.
Specifically, we construct three categories of training samples under a unified formulation that transforms an input condition into a correct (golden) RTL implementation $C^{\star}$. The input condition may consist of a specification alone or a specification paired with an existing design that must be refined or corrected. We denote the input as $X$ and define the task as: $X\rightarrow C^{\star}$, where $X$ takes different forms:
$X=\begin{cases}S&\text{(Spec-to-RTL)}\cr(S,C_{\text{base}})&\text{(Editing)}\cr(S,C_{\text{bug}})&\text{(Debugging)}\end{cases}$ (1)
Here, $S$ denotes the specification, which includes functional requirements and, depending on the task, may additionally include feature requests or diagnostic information such as signal mismatches. $C_{\text{base}}$ represents a partial implementation missing required functionality, while $C_{\text{bug}}$ denotes a faulty implementation containing design errors. Both editing and debugging tasks therefore require the model to reason over existing code context and perform targeted transformations, rather than synthesizing designs from scratch. In the following, we describe the steps of our SDG pipeline in detail.
High-Quality RTL Seed Collection
We begin by collecting approximately five million RTL scripts from public repositories and open hardware projects. To obtain a clean pool of seed designs, we apply deduplication, remove machine-generated artifacts (e.g., netlists), filter extreme outliers (30–2,000 lines per module), and validate syntax using Icarus Verilog (iverilog) [31]. We further perform benchmark decontamination using the Jaccard similarity metric [10], discarding any sample with similarity greater than $0.8$ to evaluation benchmarks. After these stages, we retain 157K high-quality RTL scripts, which we use as seed implementations to prompt LLMs for synthesizing diverse RTL problems and corresponding golden solutions.
Specification-to-RTL Data Construction
Unlike prior works that rely on chain-of-thought (CoT) reasoning supervision, we intentionally avoid training on CoT traces, since they substantially increase inference latency when the model is deployed within an agentic loop. Instead, we use a small set of human-crafted specification-to-RTL examples only as in-context demonstrations to teach frontier LLMs the desired task format and complexity. The large-scale synthetic corpus is then constructed from the collected RTL seeds in Step (a): for each seed RTL implementation, we prompt frontier LLMs, including GPT-OSS-120B and DeepSeek-R1, to generate diverse synthetic specification–code pairs conditioned on the seed design.
Editing and Debugging Data Construction
Beyond specification-to-RTL generation, we further construct editing and debugging samples to model realistic refinement workflows. Starting from the collected RTL seed implementations, which are treated as golden designs for editing and debugging tasks, LLMs are prompted to generate intermediate variants, including simplified versions with missing functionality and buggy versions with injected design errors. Simplified variants are created by removing advanced features (e.g., pipeline stages), replacing optimized control logic with baseline implementations, or restricting corner-case handling such as overflow detection or boundary conditions. Buggy variants are produced by introducing realistic RTL faults, including incorrect state transitions in finite-state machines, timing or handshake protocol violations, off-by-one counter logic, and related design mistakes.
Given these intermediate implementations, LLMs then generate corresponding specifications that either describe the required feature extensions (for editing) or provide diagnostic information (for debugging). Diagnostic specifications are constructed by analyzing the injected faults in the buggy implementation and generating failure descriptions consistent with the underlying errors. The resulting training instances require recovering the correct implementation from incomplete or faulty inputs, encouraging the model to learn iterative refinement and error-correction behaviors.
Filtering and Quality Control
All generated samples undergo multi-stage validation to ensure correctness and diversity. We first perform syntax validation and benchmark decontamination to ensure structural correctness and prevent data leakage [10] [33]. We then apply semantic alignment scoring using an LLM-as-Judge framework [11], which evaluates the consistency between specifications and implementations on a 1–5 scale. To construct the evaluation framework, we first apply the judge to a subset of training samples and manually verify whether the assigned scores are reasonable. If inconsistencies are observed, we refine the evaluation prompt and repeat the process until the scoring behavior is satisfactory. The finalized judge is then applied to the full dataset, and samples with scores below 3 are discarded to maintain data quality. This strategy provides an efficient way for large-scale SDG while maintaining sufficient quality control. Through this pipeline, we curate approximately 1.7 million high-quality training instances spanning diverse RTL coding problems.